# Decoding Quantum LDPC Codes using Collaborative Check Node Removal

### v1

@@ III-A Trapping set analysis
For QLDPC codes there exists two class of harmful TS configurations: Classical- type trapping sets (CTS) and Quantum trapping sets (QTS). The definitions of CTS and QTS are briefly covered in III.1 and III.2, which we adopt from [11].
@@
A classical-type trapping set is a set of qubits which either do not converge or they are adjacent to unsatisfied stabilizer checks after a predetermined number of iterations of a syndrome based iterative decoder.
For example in Fig. 1(a), we show a $(3,3)$ CTS obtained from the $\mathbf{H}_{Z}$ of the same $[[882,24]]$ GHP code mentioned in (11). As can be seen from the qubit and check node indexing, the CTS is formed only due to the contributions of a particular circulant matrix, namely $b^{T}I_{n}$. This type of CTS formation is also discussed in [18]. The min-sum iterative decoder in II-D fails to converge in this scenario due to the symmetric message passing rules. For example, for a received syndrome $\mathrm{s}=\{c_{0},c_{1},c_{2},c_{6},c_{7},c_{12}\}$, the min-sum BP based decoder oscillates between an all-zero error pattern and an error pattern of $\{v_{0},v_{1},v_{6}\}$, as shown in Fig 2.
@@
The other type of TS that often occurs in the QLDPC codes are the Quantum Trapping sets (QTS) as defined in III.2. These trapping sets do not have any odd-degree check nodes. Thus, a QTS with $a$ number of data qubits is referred as a $(a,0)$ trapping set. Fig. 1(b) shows a typical QTS formed for the same GHP code with $6$ data qubits in $\mathcal{T}_{s}$.
@@
The two disjoint symmetric subsets of the $(6,0)$ TS, can be identified as $V_{a}=\{v_{0},v_{351},v_{405}\}$ and $V_{b}=\{v_{477},v_{478},v_{483}\}$. Further, as per the definition III.2. $V_{a}$ and $V_{b}$ have the same set of odd-degree stabilizer check neighborhood. That is, $N(V_{a})=N(V_{b})=\{c_{0},c_{1},c_{6},c_{405},c_{406},c_{411},c_{351},c_{352},c_{367}\}$. Following the Lemma 1 from [11], error patterns of cardinality less or equals $3$ on any of the subset is a harmful configuration for the $(6,0)$ TS. For instance any $e_{a}\subseteq V_{a}:|e_{a}|\leq|V_{a}|$ has an exact twin $e_{b}\subseteq V_{b}$. This results into an oscillation between $e_{a}\oplus e_{b}$ and an all zero error at each iterations of the min-sum decoder discussed in II-D.
@@
Note that, unlike CTS, the common stabilizer neighborhood of the $(6,0)$ QTS is itself the set of odd degree check nodes for each of the disjoint subsets and therefore emerges as the unsatisfied check nodes satisfying $\mathtt{UNSAT}\subseteq N(V_{a})=N(V_{b})$. This requires different measures to approach how the iterative message passing can be improved.
III-B Use of qubit “separation" to improve iterative decoders
We now show that the iterative decoder can leverage the increased “separation” of the trapped data qubits. We observe that increasing the separation of data qubits has a direct implication towards the decoder’s performance. The oscillatory behavior of the TS discussed in the previous section is majorly contributed by the mis-satisfied check nodes of the TS. For example the $(3,3)$ TS in Fig 1(a)- $c_{1}$, $c_{6}$ are the mis-satisfied check nodes for trapped $v_{0}$. The term mis-satisfied justifies the dilemma they produce as they although are satisfied, passes mis-information, resulting in a non-convergence scenario for $v_{0}$. This phenomenon is also reflected in Fig 2. This leads to the intuition behind the separation of a qubit. It is essentially a measure of how far an incorrect trapped qubit is from another trapped qubit in the computation tree [19]. The core notion behind this measure is to identify the mis-satisfied check nodes vulnerable to decoding.
The computation tree (CT) is often used to analyze the performance of iterative decoders in classical settings [20, 19]. Consider a trapping set $\mathcal{T}_{s}$, with a set of odd degree data qubits $V_{1}$. It is built either from a trapped qubit $v\in V_{1}$ or from an unsatisfied check node $c\in C_{1}$ as a root. From the root, we add descendant check and data qubits according to the iterative message passing (or, the Tanner graph of the code). We denote a CT for $\mathcal{K}$ iterations of complete message passing cycles (i.e., qubit $\rightarrow$ check $\rightarrow$ qubit), with a qubit $v$ as root by $T_{\mathcal{K}}(v)$, and if the root is an unsatisfied check node $c$, then it is denoted by $T_{\mathcal{K}}(c)$. Therefore, $\mathcal{K}$ also indicates the level of the computation tree and from here onward we will alternately use it to denote iteration of the min-sum decoder as well as the level of a computation tree. Based on the computation tree of the trapped qubits for both CTS and QTS, the separation of the same is calculated, described in definition III.3 and III.4.
Definition III.3 (Qubit separation for CTS).
@@ III-B Use of qubit “separation" to improve iterative decoders
In Fig. 5 of the appendix, we show the computation tree $T_{4}(v_{0})$ for the $(3,3)$ CTS. We can see that the qubit $v_{0}$ has a separation of $\mathcal{K}=2$, as at $\mathcal{K}=3$ there exists $v_{6}\in V_{1}$. Also, if there are $a$ qubits in $\mathcal{T}_{s}$, then separation of each of those qubits represented in an one dimensional vector of length $a$ is called the separation vector of that $\mathcal{T}_{s}$. Note at level $\mathcal{K}=3$ of $T_{4}(v_{0})$, if we remove $c_{6}$, the separation of $v_{0}$ is increased to $3$, i.e. $\mathcal{K}=3$ for $v_{0}$. Thus, an increase in separation is intrinsic to the removal of mis-satisfied check nodes. This in-turn helps provide good information to the iterative decoders, see appendix VI-A for details. For QTS, we propose a slightly different definition for the qubit separation.
Definition III.4 (Qubit separation for QTS).
@@
If a QTS $\mathcal{T}_{s}$ has degenerate subsets $V_{a}$ and $V_{b}$, then for an erroneous data qubit node $v\in V_{a}$ with atleast one degree-one check node $c$, is said to be $\mathcal{K}$ separated, if within $\mathcal{K}$ number of message passing iterations of the decoder, there exists no more data qubit $u\in V_{a}$ as a descendant of $c$ in $T_{\mathcal{K}}(c)$.
Similarly, we define the separation for the qubits in $V_{b}$. We note that for QTS, the contributions of the harmful information are channelled through the check nodes connecting the disjoint symmetric subsets. These check nodes play the same role as the mis-satisfied check nodes of the CTS. Although they themselves can be the only unsatisfied check nodes for the harmful error configurations, as we discussed before. In Appendix VI-A, we show for $(6,0)$ QTS, removal of check node pairs from $T_{k}(v_{i})$ (in accordance with definition III.4) has an impact in increasing the separation for $v_{i}$, where $v_{i},v_{j}\in\{V_{a},V_{b}\}$. Further, we provide evidences in appendix VI-A, showing that removal of these checks help the iterative decoder predict correct syndromes, which was not possible through the iterative decoder under normal consitions.
In the following section, we first draw an analogy with the classical results from Zhang et al. [21] to show that increased separation of a trapped node successfully breaks typical unsuccessful decoding scenarios of the iterative decoder. Further, we perform numerical memory experiments to:
@@
    show the advantage of removing mis-satisfied check nodes through a collaborative decoding approach,
@@ IV Decoding via Collaborative Check Node Removal
As discussed in the previous section, removing the correct check node from the computation tree of a trap node increases its separation and improves the performance of iterative decoders. However, knowing the exact level in the CT from where the check node removal can be done requires the TS information. For example, Fig. 5 in Appendix VI-A shows clearly that the candidate check nodes for removal are in level $\mathcal{K}=3$ of the CT $T_{4}(v_{0})$. These candidate check nodes are the same mis-satisfied check nodes, information of which can only be known from the TS. Thus, a natural practical way to approach this scenario is through the same probabilistic approach of [13], which increases the separation of the trapped nodes without knowing the particulars of the TS of the quantum code. From here onward, we denote neighbor of a qubit and check as $N(v)$, $N(c)$ respectively and alternately use $t$ and $\mathcal{K}$ for algorithmic syntax purposes.
IV-A Weak check node removal

### v2

@@ III-A Trapping set analysis
For QLDPC codes, there exist two classes of harmful TS configurations: Classical-type trapping sets (CTS) and Quantum trapping sets (QTS). The definitions of CTS and QTS are briefly covered in III.1 and III.2, which we adopt from [12].
@@
A classical-type trapping set is a set of qubits that either do not converge or are adjacent to unsatisfied stabilizer checks after a predetermined number of iterations of a syndrome-based iterative decoder.
For example in Fig. 1(a), we show a $(3,3)$ CTS obtained from the $\mathbf{H}_{Z}$ of the $[[882,24]]$ GHP code discussed in (11). As can be seen from the qubit and check node indexing, the CTS is formed only due to the contributions of a particular circulant matrix of the form $b^{T}I_{n}$. This type of CTS formation is also discussed in [21]. The min-sum iterative decoder in II-D fails to converge in this scenario due to the symmetric message passing rules. For example, for a received syndrome $\mathrm{supp}(\mathrm{s})=\{c_{0},c_{1},c_{2},c_{6},c_{7},c_{12}\}$, indicating the violated stabilizer checks, the min-sum BP based decoder oscillates between an all-zero error pattern and an error pattern of $\text{supp}(e)=\{v_{0},v_{1},v_{6}\}$, as shown in Fig 2.
@@
The other type of TS that often occurs in the QLDPC codes is the Quantum Trapping sets (QTS) as defined in III.2. These trapping sets do not have any odd-degree check nodes. Thus, a QTS with $a$ number of data qubits is referred as a $(a,0)$ trapping set. Fig. 1(b) shows a typical QTS formed for the same GHP code with $6$ data qubits in the $\mathcal{T}_{s}$.
@@
The two disjoint symmetric qubit subsets of the $(6,0)$ TS, can be identified as $V_{a}=\{v_{0},v_{351},v_{405}\}$ and $V_{b}=\{v_{477},v_{478},v_{483}\}$. Further, as per the definition III.2. $V_{a}$ and $V_{b}$ have the same set of odd-degree stabilizer check neighborhood. That is, $N(V_{a})=N(V_{b})=\{c_{0},c_{1},c_{6},c_{405},c_{406},c_{411},c_{351},c_{352},c_{367}\}$. Following the Lemma 1 from [12], error patterns of cardinality less or equals $3$ on any of the subset is a harmful configuration for the $(6,0)$ TS. For instance any $e_{a}\subseteq V_{a}:|e_{a}|\leq|V_{a}|$ has an exact twin $e_{b}\subseteq V_{b}$. This results into an oscillation between $e_{a}\oplus e_{b}$ and an all zero error at each iterations of the min-sum decoder discussed in II-D.
@@
Note that, the common stabilizer check neighborhood of the $(6,0)$ QTS is itself the set of odd degree check nodes, when we consider the disjoint subsets independently. Therefore, the potential violation of check nodes occur from $N(V_{a})$ or, $N(V_{b})$ (i.e. $\mathtt{UNSAT}\subseteq N(V_{a})=N(V_{b})$) for any of the possible harmful error patterns for the min-sum decoder. This particular phenomena requires different measures compared to the approach for CTS, while attempting to improve the iterative decoder’s performance using qubit separation, which we elaborate next.
III-B Mitigate TS configurations through “qubit separation"
We now show that the iterative decoders can leverage from the increased “separation” of the trapped data qubits. The oscillatory behavior of the TS discussed in the previous section is majorly contributed by the mis-satisfied check nodes of the TS, as discussed in [16]. In the context of CTS for quantum codes, one can assume that the mis-satisfied nodes are the satisfied stabilizer checks, which are adjacent to an even number of incorrect qubits. Therefore, one can relate that these types of nodes are specific to the CTS. In a $(a,b)$ CTS, mis-satisfied check nodes can be present in any amount. For example, in case of the $(3,3)$ TS in Fig 1(a); $c_{1}$, $c_{6}$ and $c_{7}$ are the mis-satisfied check nodes. The term mis-satisfied justifies the dilemma of passing incorrect messages during decoding iterations, although being satisfied. For instance the messages due to $c_{1}$ and $c_{6}$, results in a non-convergence for $v_{0}$. This phenomenon is also reflected in Fig 2. This leads us to a natural intuition, where if we can identify such mis-satisfied stabilizer check nodes, removal of such can increase the possibility of the iterative decoder’s convergence, through a message of correct information.
Kang et al. proposed the use of bit separation for the classical TS [16]. It is essentially a measure of how far an incorrect trapped bit is from another bit trapped in the same TS, found along the computation tree of the former bit [22]. The computation tree (CT) is often used to analyze the performance of iterative decoders in classical settings [22, 23]. We adopt the definition of the CT from [16, 22] and in the context of quantum codes, it is as follows:
Definition III.3 (Computation Tree).
The computation tree $\mathcal{T}_{\mathcal{K}}(t)$ for an iterative decoder on a code is constructed by picking a root node ‘$t$’ corresponding to either a qubit or a stabilizer check node and then iteratively adding ‘$\mathcal{K}$’ levels of edges and leaf nodes according to each complete iterations (i.e., qubit (check) $\rightarrow$ check (qubit) $\rightarrow$ qubit (check)) of the messages passed for the iterative decoder.
We can now explore the role of CT construction in the analysis of qubit separation. Consider a TS $\mathcal{T}_{s}$ having a set of odd-degree data qubits and stabilizer checks $V_{1}$ and $C_{!}$ respectively. We start building the CT from a trapped qubit $v\in V_{1}$ or from an unsatisfied check node $c\in C_{1}$ as a root. From the root, we add descendant stabilizer check or data qubits according to the iterative message passing, which directly follows from the Tanner graph of the code. Due to this, for the rest of the article we use $\mathcal{K}$ to denote the level of CT as well as the number of complete iterations of iterative decoders on similar note. We now define the separation of a qubit trapped inside a CTS:
Definition III.4 (Qubit separation for CTS).
@@ III-B Mitigate TS configurations through “qubit separation"
$v_{0}$$c_{0}$$c_{1}$$c_{6}$$v_{477}$$v_{62}$$v_{57}$$v_{513}$$v_{567}$$c_{5}$$c_{62}$$v_{5}$$v_{4}$$c_{6}$$c_{11}$$v_{6}$$v_{10}$$v_{11}$$v_{56}$$c_{57}$$v_{57}$123
Figure 3: Computation tree of the trapped qubit $v_{0}$ of CTS $(3,3)$. We show one of the trapped qubits as a descendant of $v_{0}$ at layer $\mathcal{K}=3$. At each level, the nodes (data or stabilizer check) are added according to where the messages are being passed from parent to child based on the iterative decoder.
In Fig. 3, we show the computation tree $T_{4}(v_{0})$ for the $(3,3)$ CTS of $[[882,24]]$ GHP code. We observe that the qubit $v_{0}$ has a separation of $\mathcal{K}=2$, as at $\mathcal{K}=3$ there exists $v_{6}\in V_{1}$. From Fig. 3, we also observe at level $\mathcal{K}=3$ of $T_{4}(v_{0})$, if we remove $c_{6}$, the separation of $v_{0}$ is increased. We also note that $c_{6}$ is indeed a mis-satisfied node as discussed above. Therefore, an increase in separation directly relates to the removal of mis-satisfied check nodes. Proposition $2$ of [16] or theorem 1 of [24], supports that a sufficiently large separation leads to correctable root trapped qubit of the CT by an iterative decoder. The same can be observed from Table. I.
Removed Stabilizer check nodes Affected trapped qubit New syndrome bit predictions by min-sum decoder
$c_{6}$ $v_{0}$, $v_{6}$ $[c_{0},c_{1},c_{7},c_{12}]$
$c_{1}$ $v_{0}$, $v_{1}$ $[c_{0},c_{2},c_{6},c_{7}]$
$c_{7}$ $v_{1}$, $v_{6}$ $[c_{1},c_{2},c_{6},c_{12}]$
TABLE I: Direct check node removal showing the positive impact on the prediction of erroneous trapped qubits by min-sum BP decoder for a $(3,3)$ CTS.
The nature and harmful effect of QTS is intrinsically different than the CTS. Considering the differences, we propose a slightly different definition for the qubit separation in the context of QTS.
Definition III.5 (Qubit separation for QTS).
@@
If a QTS has degenerate subsets $V_{a}$ and $V_{b}$, then for an erroneous data qubit node $v\in V_{a}$ with atleast one degree-one (w.r.t. the same subset) check node $c$, is said to be $\mathcal{K}$ separated, if within $\mathcal{K}$ number of message passing iterations of the decoder, there exists no more data qubit $u\in V_{a}$ as a descendant of $c$ in $T_{\mathcal{K}}(c)$.
The same definition extends towards the separation of qubits in $V_{b}$. In the context of QTS, the mis-satisfied check nodes do not exist. Instead, a subset of the common stabilizer check neighborhood is responsible for passing incorrect information to the erroneous data qubits in the QTS. These check nodes are present in level $1$ of the computation tree with the erroneous data qubit as it’s root. Removal of these subset of nodes improves the separation of the root erroneous qubit and helps iterative decoder pass correct information to the erroneous qubit.
Lemma 2.
A QLDPC code with data qubits having degree $d_{v}$, has a total of $d_{v}(d_{v}-1)$ number of stabilizer check nodes $c_{r}:c_{r}\in N(V_{a})=N(V_{b})$, which limits the separation of any erroneous data qubit $v_{r}\in V_{a}$ or, $v_{r}\in V_{b}$ with $V_{a}$ and $V_{b}$ together forming the QTS.
Proof.
The proof of the above is very simple and is a direct consequence of the isomorphic property of the QTS. As in a typical QTS, there are no odd-degree stabilizer check nodes; therefore, all the qubits in the QTS exhibit $d_{v}$ number of adjacent stabilizer check nodes. Further from the definition of QTS, we can conjecture that the isomorphic nature of the two disjoint subsets results in a set of qubits at level $2$ of $T(v_{r})\forall v_{r}\in V_{i}$, which limits the qubit separation. We assume these qubits are child nodes of the parent $p$ number of check nodes. Each leaf data qubit nodes at level $1$ of $T(v_{r})$ has $d_{v}-1$ number of child check nodes at level $2$, therefore each of these $d_{v}-1$ number of nodes has one child data qubit in the $T(v_{r})$, which contributes to limiting qubit separation. So, $p=d_{v}-1$. Now finally we need to look at how many parent data qubit nodes at level $1$ of the CT, has these $d_{v}-1$ number of check nodes at level $2$. This number is simply $d_{v}$, because the root qubit has $d_{v}$ number of child check nodes and from each such check nodes one of the descendant qubit in level $1$ belongs to the other isomorphic qubit subset of the QTS. Therefore, in total $d_{v}(d_{v}-1)$ number of check nodes are connected to qubits at level $2$, which contributes to the limiting condition for qubit separation. ∎
Lemma 3.
Any erroneous qubit $v_{r}\in V_{i}$, where $V_{i}$ are the isomorphic qubit subsets of the QTS; can be corrected by min-sum BP decoder if the $d_{v}(d_{v}-1)$ number of check nodes from the level $1$ of the CT, $T(v_{r})$, which are limiting the separation of $v_{r}$; are removed from the code’s Tanner graph.
Proof.
To proof the above lemma, we analyze the message received by one of the erroneous data qubits in the QTS. For instance assume $v_{r}\in V_{i}$, where $V_{i}$ is one of the degenerate qubit subsets of the QTS and $c_{r}\in N(V_{i})$ is one of the common stabilizer check nodes, which connects two different disjoint qubit subsets of the QTS. Also, $T(v_{r})$ is the computation tree with $v_{r}$ as it’s root. We now observe the flow of information in $T(v_{r})$ from leaf nodes towards the root node. For the min-sum BP decoder to be able to correct erroneous $v_{r}$, correct information must be passed from $c_{r}$ to root $v_{r}$ in $T(v_{r})$. For this to be done, all the information passed to $c_{r}$ from it’s descendants must be correct. The correct passage of information must satisfy the following equality for messages passed from any child node $v_{i}$ to the parent $c_{r}$,
$\mathrm{sign}(\lambda_{i})=\mathrm{sign}(m_{v_{i}\rightarrow c_{r}});\forall v_{i}\in N(c_{r})\backslash v_{r}.$ (20)
Eq. (20) imply that as information passed into $c_{r}$ is correct, sign of all the messages passed into $c_{r}$ should be same as that of the message passed into data qubit nodes $v_{i}$ from the channel. Now, assuming phenomenological bit flip noise channel, the magnitude of the messages received by the data qubit nodes from the channel are equal. Therefore for $d_{v}$ regular QLDPC code, the messages send from any data qubit $v_{i}$ in the same level of $T(v_{r})$ has equal magnitude and follows
$|m_{v_{i}\rightarrow c_{j}}|$ $=|\lambda_{i}|+\sum_{c_{j^{\prime}}:N(v_{i})\backslash c_{j}}|m_{c_{j^{\prime}}\rightarrow v_{i}}|.$ (21)
Therefore, if we consider the effect of the stabilizer check nodes at level $2$ of $T(v_{r})$, we directly observe violation of Eq. (20). This is because of the following condition
$\mathrm{sign}(m_{c_{j}\rightarrow v_{i}})\neq\mathrm{sign}(\lambda_{i}),$ (22)
where if $v_{r}\in V_{i}$, then $v_{i}\in V_{j}$; such that $V_{i}\cap V_{j}=\emptyset$. At the junction of level $1$ and level $2$, qubit $v_{i}$ is adjacent to $d_{v}-1$ child check nodes and all such nodes follow Eq. (22). Therefore, the check nodes at level $2$ of $T(v_{r})$, which limit the separation of $v_{r}$, also pass incorrect information to check node $c_{r}$ adjacent to $v_{r}$, which leads to net incorrect information passed to $v_{r}$ by the min-sum BP decoder. Therefore, removal of all such $d_{v}(d_{v}-1)$ number of check nodes from level $2$ of $T(v_{r})$ ensures correct information passed to the erroneous qubits and, as a result, ensures correctness of the erroneous qubits in the QTS. ∎
$v_{0}$$c_{0}$$c_{1}$$c_{6}$$v_{478}$$v_{477}$$v_{483}$$c_{406}$$c_{352}$$v_{405}$$v_{351}$$c_{351}$$c_{405}$$v_{405}$$v_{351}$$c_{411}$$c_{357}$$v_{405}$$v_{351}$12
$v_{405}$$c_{405}$$c_{406}$$c_{411}$$v_{478}$$v_{477}$$v_{483}$$c_{352}$$c_{1}$$c_{357}$$c_{6}$$v_{0}$$v_{351}$$v_{351}$$v_{0}$$c_{0}$$c_{351}$$v_{0}$$v_{351}$12
$v_{351}$$c_{357}$$c_{352}$$c_{351}$$v_{478}$$v_{483}$$v_{477}$$c_{0}$$c_{405}$$v_{0}$$v_{405}$$c_{6}$$c_{411}$$v_{0}$$v_{405}$$c_{1}$$c_{406}$$v_{0}$$v_{405}$12
Figure 4: Computation tree and the potential structure indicating the scope of separation improvement for all the trapped qubits of one of the disjoint subsets of the $(6,0)$ QTS of $[[882,24]]$ GHP code. Note the dotted branches indicated in each of the computation trees. These are the branches bringing flexibility towards separation increments in a QTS enabled by definition III.2. The square check adjacent to leaf blue nodes are the candidate set for removal. The two propositions discussed above can be verified from the diagram.
In Fig. 4, we show the CT with root qubits from one of the degenerate qubit subsets of the $(6,0)$ QTS of the $[[882,24]]$ GHP code. This code belongs to a family of GHP codes with $d_{v}=3$. We validate the results of Lemma 2 and observe that for each of the erroneous trapped qubits, there are exactly $d_{v}(d_{v}-1)=6$ stabilizer check nodes at level $2$ of the computation tree, which are limiting the separation of the trapped qubits.
Removed Stabilizer check nodes Affected trapped qubit New syndrome bit predictions by min-sum decoder
$[c_{406},c_{352},c_{351},c_{405},c_{357},c_{411}]$ $v_{0}$ $[c_{0},c_{1},c_{6}]$
$[c_{0},c_{405},c_{1},c_{406},c_{6},c_{411}]$ $v_{351}$ $[c_{351},c_{352},c_{357}]$
$[c_{0},c_{351},c_{1},c_{352},c_{6},c_{357}]$ $v_{405}$ $[c_{405},c_{406},c_{411}]$
TABLE II: Direct check node removal showing the positive impact on the prediction of erroneous trapped qubits by min-sum BP decoder for a $(6,0)$ QTS.
In table II, we validate the results of lemma 3 and observe that removal of those separation limiting $6$ stabilizer check nodes actively improves the min-sum BP decoder’s performance and ensures correction of the root erroneous trapped qubit.
In the following section, we first discuss an analogy with the classical results from Zhang et al. [24] to discuss the limitations imposed by the probabilistic decoding algorithm. We then use the concepts of “information measurement" from [25] to improve the possibility of an accurate set of stabilizer check removals from the code’s Tanner graph and discuss a new algorithm which provides a better success rate than the min-sum BP algorithm. We perform numerical memory experiments to:
@@
    show the advantage of removing proper check nodes through a collaborative decoding approach,
@@ IV Decoding via Collaborative Check Node Removal
As discussed in the previous section, removing the check nodes from the computation tree of an erroneous trapped node increases its separation and improves the performance of min-sum BP decoder. However, knowing the exact level in the CT from where the check node removal can be done requires the TS information, specially for the CTS. Even so, identification of such erroneous trapped qubits also requires the information of TS. For example, Fig. 3 and 4 in Section III-B clearly identify the removable candidate check nodes from the CT of the erroneous trapped qubits. These removable check nodes are present at level $\mathcal{K}=3$ and $\mathcal{K}=2$ of the CT of respective erroneous qubits belonging to the $(3,3)$ CTS and $(6,0)$ QTS, respectively. These candidate check nodes are a set of mis-satisfied check nodes and a subset of common stabilizer check nodes of the CTS and QTS, respectively, information of which can only be available upon knowing about the corresponding TS. Therefore, a natural, practical way to approach this scenario is through the same probabilistic approach of [16], which increases the separation of the trapped nodes without knowing the particulars of the TS of the quantum code. Next, we first discuss the method proposed by Kang et al. in the context of QLDPC codes. We discuss the concept of weak check nodes and how it allows us to derive an algorithm that can improve the min-sum decoder’s performance without knowing the particulars of the TS. Following it, we propose an improved algorithm, namely Quantum Collaborative Check Node Removal (QCCNR); which has more potential to eliminate the vulnerable stabilizer check nodes from the quantum code and achieve better success rates using a collaborative architecture.
IV-A Weak stabilizer check node removal